Artisan Works is owned and operated by a craftsman who makes replicas of historic firearms for​ museums, sportsmen, and collectors. The data are as​ follows:

Sales price per unit

​$780

Variable cost per unit

470

Fixed costs per month

​8,370

If Artisan expects to sell 60 units per​ month, what is his margin of safety expressed in units per​ month?
Answer #1

Contribution margin=Sales-Variable cost  

=(780-470)=$310 per unit

Breakeven=Fixed expenses/Contribution margin

=(8370/310)=27 units

Margin of safety=Total sales-Breakeven sales

=(60-27)

=33 units